⦁ Ultraviolet light of wavelength 207 nm causes photoemission from a surface? The stopping potential is 2 V. Find (2) the work function in eV; (b) the maximum speed of the photoelectrons
1
Expert's answer
2020-05-07T08:41:31-0400

(a) the work function

"\varphi=\frac{hc}{\lambda}-eV\\
=\frac{1240\:\rm eV\cdot nm}{207\:\rm nm}-2\:\rm eV=4\:\rm eV."

(b) the maximum speed of the photo-electrons

"v=\sqrt{\frac{2eV}{m}}=\sqrt{\frac{2\times 1.6\times 10^{-19}\times 2}{9.1\times 10^{-31}}}=8.4\times 10^5 \:\rm m/s."